i cant figure out how to put this in a algebra equation. "At 1:30p.m., Tom leaves in his boat from a dock and heads south. He travels at a rate of 25 miles per hour. Ten minutes later, Mary leaves the same dock in her speedboat and heads after Tom. If she travels at a rate of 30 miles per hour, when will she catch up to Tom?
Answer by richwmiller(17219)

25(x+10/60)=30x
25x+10/60*25=30x
250/60=5x
50/60=x
add 50 minutes to 1:40 pm
1:90 pm=2:30 pm